OpenStudy (anonymous):

The probability that Shania is on time for school is . Find the probability that Shania arrives on time for school for the next 5 days. Express your answer as a percent, to the nearest tenth of a percent.

OpenStudy (anonymous):

P(Shania arrives on time for school for the next 5 days) = P(arrived on time on first day) x P(arrived on time on second day) x P(arrived on time on third day) x P(arrived on time on fourth day) x P(arrived on time on fifth day) = (1/2) x (1/2) x (1/2) x (1/2) x (1/2) = (1/2)^5 = ?

OpenStudy (mathmate):

p=probability that she arrives on time for one day. for two days, it's p^2 ...

OpenStudy (anonymous):

Out of 1,100 discs tested, 13 are defective. Estimate the number of defective discs in a batch of 41,000.

OpenStudy (mathmate):

I would just use the expected value, such as np.